OpenStudy (anonymous):

Find the intercepts and asymptotes. t(x) =(x^3 − 8x^2)/(x^3 − 3x − 2)

OpenStudy (anonymous):

\[(x^2(x-8))/((x-2)(x+1)^2)\] zeros is when top equal zero so when \[x^2(x-8)=0\] x=0 and x=8 vertical asymptote when bottom equal zero sowhen : \[(x-2)(x+1)^2=0\] x=2 and x=-1 horizontal asymptote since the highest power is the same its the quotient of the coefficients 1/1=1